a boy throws a stone horizentally off a cliff.it hits the sea 2 seconds later,at a distance of 40 meters from the foot of the cliff.                                                                                             calculate the                                                                             (a)height of the cliff                                                                   (b) intial speed of the stone                                                      © direction in which the stone is moving when it strikes the water.

Grade:12th pass

1 Answers

Arun
4 years ago
Height = 0.5 g t*t= 0.5*10*2*2= 20Initial speed = horizontal dist/ time= 40/ 2 = 20 m/s scalar quantityDirection of stone is in the direction of initial speed if no obstacle is present